OpenStudy (anonymous):

What kind of function best models the data? X: Y: 0 30 1 15 2 7.5 3 3.75 4 1.875

OpenStudy (anonymous):

I need help D:

OpenStudy (amistre64):

it seems to be halfing itself

OpenStudy (amistre64):

y = (1/2)^x

OpenStudy (amistre64):

\(y = (1/2)^x\)

OpenStudy (amistre64):

+29 ?

OpenStudy (amistre64):

+30 perhps lol

OpenStudy (anonymous):

do you know what kind of equation it is though?

OpenStudy (amistre64):

its a recursice equation

OpenStudy (amistre64):

recursive

OpenStudy (amistre64):

A{0} = 30 ; A{n} = A{n-1} * (1/2)

OpenStudy (amistre64):

\[A_n = (1/2)A_{n-1}\] \[A_n = (1/2)^2 A_{n-2}\] \[A_n = (1/2)^3 A_{n-3}\]

OpenStudy (amistre64):

\[A_n = (1/2)^n A(0)\] \[A_n = (1/2)^n * 30\]

OpenStudy (anonymous):

Thanks

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=30*%281%2F2%29^n to see a graph ;)

OpenStudy (anonymous):

That helps a lot thanks :D